OpenStudy (anonymous):

the second term and the fifth term in a geometric progression is 6 and 162 respectively. Find the first term and the common ratio of the progression. please help me to solve this.

OpenStudy (anonymous):

\[a_2 = 6\] \[a_5 = 162\] \[a_2 = a*q\] \[a_5 = a*q^4\] \[a_5/a_2 = a*q^4/a*q\] \[162/6 = q^3\] \[27= q^3\] \[9= q\] \[a_2 = a*q\] \[6 = a*9\] a =?

OpenStudy (anonymous):

ok thnks i got it ^_^

OpenStudy (anonymous):

Anytime :)